Walt made an extra $9000 last year from a part-time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770 in interest. How much was invested at 8%.
I believe the answer to be $4000 invested at 8% because
$4000x.08=$320 and $5000x.09=$450 so $320+$450=$770. I Cannot however figure out how to put this in a linear equation. Please help.
: Walt made an extra $9000 last year from a part-time job. He invested part of the money at 9% and the rest at 8%. He made a total of $770 in interest. How much was invested at 8%.
I believe the answer to be $4000 invested at 8% because
$4000x.08=$320 and $5000x.09=$450 so $320+$450=$770. I Cannot however figure out how to put this in a linear equation. Please help.

Answer by checkley77(3624)

.09x+.08(9000-x)=770
.09x+720-.08x=770
.01x=770-720
.01x=50
x=50/.01
x=$5,000 amount invested @ 9%
9000-5000=$4,000 amount invested @ 8%
proof:
.09*5000+.08*4000=770
450+320=770
770=770